OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the given point. y=x^4+2x^2-x, (1,2)

OpenStudy (dumbcow):

here a good way of looking at tangent line given f(x) and point (x1,y1) \[y - y_1 = f'(x_1)(x - x_1)\]

OpenStudy (whpalmer4):

Take the first derivative to get a function for the slope of the tangent line at a point. Then use the value of the first derivative at the point to get the slope, and find the equation of the tangent line with the point-slope formula \[y-y_0 = m(x-x_0)\]

OpenStudy (anonymous):

Thanks!

OpenStudy (whpalmer4):

Should look like this when you're all done. What did you get for your answer?

OpenStudy (anonymous):

Yup, looks something similar!